-10x-7y=24
-9x-9y=0

Answer by checkley79(3341)

-10x-7y=24
-9x-9y=0
-9X=9Y
X=9Y/-9
X=-Y ANS.
-10*-Y-7Y=24
10Y-7Y=24
3Y=24
Y=24/3
Y=8 ANS.
PROOFS:
-9(-Y)-9Y=0
9Y-9Y=0
0=0
-10(-8)-7*8=24
80-56=24
24=24